LASER SPECTROSCOPY AND DYNAMICS OF THE JET-COOLED AsH_{2} FREE RADICAL

SHENG-GUI HE, Institute of Chemistry, Chinese Academy of Sciences, Zhongguancun, Beijing 100080, P.R.C.; \underline{DENNIS J. CLOUTHIER}, Department of Chemistry, University of Kentucky, Lexington, KY 40506-0055.

The Ã^{2}A_{1} - X̃^{2}B_{1} electronic transition of the jet-cooled AsH_{2} free radical has been studied by laser-induced fluorescence (LIF), wavelength-resolved emission, and fluorescence lifetime measurements. The radical was produced by a pulsed electric discharge through a mixture of arsine (AsH_{3}) and high-pressure argon at the exit of a pulsed valve. Nine vibronic bands were identified by LIF spectroscopy in the 505-400 nm region, including a long progression in the bending mode and two bands (1_{0}^{1} and 1_{0}^{1}2_{0}^{1}) involving the excited state As-H symmetric stretch. Single vibronic level emission spectra showed similar activity in the bending and symmetric stretching frequencies of the ground state. High- resolution spectra of the 0_{0}^{0} band exhibited large spin-splittings and small, resolved arsenic hyperfine splittings, due to a substantial Fermi contact interaction in the excited state. The rotational constants obtained in the analysis gave effective molecular structures of r′′_{0} = 1.5183(1) Å, θ′′_{0} = 90.75(1)^{o} and r′_{0} = 1.4830(1) Å, θ′_{0} = 123.10(2)^{o}. The excited state fluorescence lifetimes vary dramatically with rovibronic state, from a single value of 1.4 μs to many with lifetimes less than 10 ns, behavior which we interpret as signaling the onset of a predissociative process in the excited state.